give an example for a gaseous reversible reaction for which kp and kc are equal
1
Expert's answer
2014-01-31T12:17:32-0500
The values Kc and Kp are equal when the total number of moles of gaseous substances in the both parts of the equation is same:
CO2(g) + H2(g) = CO(g) + H2O(g)
NO2(g) + SO2(g) = NO(g) + SO3(g)
C(s) + O2(g) = CO2(g)